# TwinFlow: Realizing One-step Generation on Large Models with Self-adversarial Flows

### v1

@@ 4.3 Image Generation on Dedicated Text-to-image Models
SANA-Sprint-0.6B (Chen et al., 2025b) 2 6.46 0.25 0.6B 0.76 81.5
SANA-Sprint-1.6B (Chen et al., 2025b) 2 5.68 0.24 1.6B 0.77 82.1
@@
SANA-Sprint-0.6B (Chen et al., 2025b) 1 7.22 0.21 0.6B 0.72 78.6
SANA-Sprint-1.6B (Chen et al., 2025b) 1 6.71 0.21 1.6B 0.76 80.1
@@
    1-NFE setting: The efficacy of our method is particularly pronounced in the more demanding 1-NFE inference setting. Here, our models (0.6B: 0.83, 1.6B: 0.81 on GenEval) significantly outperform other leading 1-NFE methods, such as SANA-RCGM (0.78) (Sun & Lin, 2025), SANA-Sprint (0.76) (Chen et al., 2025b), FLUX-Schnell (0.69) (Labs, 2024), and SDXL-DMD2 (0.59) (Yin et al., 2024a). Notably, our 1-NFE TwinFlow-0.6B (GenEval: 0.83) exceeds the generation quality of the 40-NFE SANA-1.5-4.8B (Xie et al., 2025a) model while offering substantially greater computational efficiency.
@@ Appendix A Related Work
Diffusion (Ho et al., 2020; Dhariwal & Nichol, 2021) and flow matching (Lipman et al., 2022) models have shown impressive performance in image generation. They progressively transport a simple noise distribution to the data distribution, either by reversing a noising process or integrating a probability-flow ODE. However, this iterative nature introduces bottleneck in inference efficiency, as generating an image requires numerous sequential evaluation steps.
@@
Various methods tend to accelerate the sampling process, such as diffusion distillation (Luhman & Luhman, 2021; Salimans & Ho, 2022) and consistency distillation (Song et al., 2023; Song & Dhariwal, 2023). They typically train a student model to approximate the ODE sampling trajectory of a frozen teacher model in fewer sampling steps. While effective at moderate NFEs, these approaches depend on a frozen teacher, and quality often drops sharply in the extreme few-step regime (< 4 NFEs). Incorporating GAN-like loss into distillation (e.g., CTM (Kim et al., 2023), ADD/LADD (Sauer et al., 2024b; a), DMD/DMD2 (Yin et al., 2024b; a)) can improve sharpness and alignment at few steps. Yet these frameworks increase training complexity and instability: they typically introduce auxiliary modules (discriminators, fake-sample score functions) and still rely on a frozen teacher, leading to higher memory overhead and sensitivity to hyperparameters. For ultra-large models, this added complexity often translates to out-of-memory failures or brittle training dynamics.
@@
The tension between speed and quality is amplified in large-scale systems. For instance, Qwen-Image-20B (Wu et al., 2025a) typically requires 100 NFEs, leading to substantial latency ($\sim$40s on a single A100 for $\times1024$ resolution). Recent works distill to cut NFEs while preserving compositionality and aesthetics: LCM-style distillation for latent models (Luo et al., 2023), large-scale text-to-image distillation pipelines (e.g., PixArt-Delta (Chen et al., 2024b), SDXL distillation (Lin et al., 2024), FLUX-schnell (Labs, 2024)), and hybrid frameworks such as SANA-Sprint (Chen et al., 2025b) that combine teacher guidance with adversarial signals. Recently, Hunyuan-Image-2.1 (Team, 2025) explore MeanFlow (Geng et al., 2025) for mid-step acceleration (16 NFEs). Nevertheless, when targeting 1-2 NFEs on $\times1024$ text-to-image with 10B-20B backbones, these pipelines face practical barriers: dependence on frozen teachers, extra discriminators or score networks, unstable adversarial training, and prohibitive memory costs that hinder straightforward scaling.
@@ B.1 Detailed Implementation on Multimodal Generative Models
For training OpenUni (Wu et al., 2025c) and Qwen-Image (Wu et al., 2025a), we used the same datasets as in our text-to-image experiments but excluded ShareGPT-4o-Image (Chen et al., 2025c). For exploration on full-parameter training on Qwen-Image, since DMD (Yin et al., 2024b) employs a regression loss, we need to pre-generate offline data using Qwen-Image. To ensure a fair comparison with established baselines, we randomly sample 50K prompts from the datasets we used, and generate images using Qwen-Image. For Qwen-Image-Edit (Wu et al., 2025a), we used only the part of editing split of the ShareGPT-4o-Image dataset.
@@ B.2 Detailed Implementation on Dedicated Text-to-image Models
Considering training datasets, we use BLIP-3o-60K (Chen et al., 2025a), Echo-4o (w/o multi-reference split) (Ye et al., 2025), and ShareGPT-4o-Image (Chen et al., 2025c). Together, these three instruction tuning datasets comprise approximately 200,000 text-to-image samples.
@@ B.3 Exploration on Image Editing
We conducted a preliminary exploration of our method on image editing tasks using a subset of approximately 15,000 square images from Chen et al. (2025c). We fine-tuned the Qwen-Image-Edit model (Wu et al., 2025a) using LoRA at a fixed $\times512$ resolution, with training configurations detailed in Table 6. Note that we use a low and fixed resolution during training; during testing, we use a resolution that is the same as the input image size. Our Qwen-Image-Edit-TwinFlow can effectively edit images in 2 to 4 NFEs. With 2 NFEs, it achieves a score of 3.47 on the ImgEdit, surpassing all multi-step models except Qwen-Image-Edit itself.

### v2

@@ 4.3 Image Generation on Dedicated Text-to-image Models
SANA-Sprint-0.6B (Chen et al., 2025c) 2 6.46 0.25 0.6B 0.76 81.5
SANA-Sprint-1.6B (Chen et al., 2025c) 2 5.68 0.24 1.6B 0.77 82.1
@@
SANA-Sprint-0.6B (Chen et al., 2025c) 1 7.22 0.21 0.6B 0.72 78.6
SANA-Sprint-1.6B (Chen et al., 2025c) 1 6.71 0.21 1.6B 0.76 80.1
@@
    1-NFE setting: The efficacy of our method is particularly pronounced in the more demanding 1-NFE inference setting. Here, our models (0.6B: 0.83, 1.6B: 0.81 on GenEval) significantly outperform other leading 1-NFE methods, such as SANA-RCGM (0.78) (Sun and Lin, 2025), SANA-Sprint (0.76) (Chen et al., 2025c), FLUX-Schnell (0.69) (Labs, 2024), and SDXL-DMD2 (0.59) (Yin et al., 2024a). Notably, our 1-NFE TwinFlow-0.6B (GenEval: 0.83) exceeds the generation quality of the 40-NFE SANA-1.5-4.8B (Xie et al., 2025a) model while offering substantially greater computational efficiency.
@@ Appendix A Related Work
Diffusion (Ho et al., 2020; Dhariwal and Nichol, 2021) and flow matching (Lipman et al., 2022) models have shown impressive performance in image generation. They progressively transport a simple noise distribution to the data distribution, either by reversing a noising process or integrating a probability-flow ODE. However, this iterative nature introduces bottleneck in inference efficiency, as generating an image requires numerous sequential evaluation steps.
@@
Various methods tend to accelerate the sampling process, such as diffusion distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022) and consistency distillation (Song et al., 2023; Song and Dhariwal, 2023). They typically train a student model to approximate the ODE sampling trajectory of a frozen teacher model in fewer sampling steps. While effective at moderate NFEs, these approaches depend on a frozen teacher, and quality often drops sharply in the extreme few-step regime (< 4 NFEs). Incorporating GAN-like loss into distillation (e.g., CTM (Kim et al., 2023), ADD/LADD (Sauer et al., 2024b; a), DMD/DMD2 (Yin et al., 2024b; a)) can improve sharpness and alignment at few steps. Yet these frameworks increase training complexity and instability: they typically introduce auxiliary modules (discriminators, fake-sample score functions) and still rely on a frozen teacher, leading to higher memory overhead and sensitivity to hyperparameters. For ultra-large models, this added complexity often translates to out-of-memory failures or brittle training dynamics.
@@
The tension between speed and quality is amplified in large-scale systems. For instance, Qwen-Image-20B (Wu et al., 2025a) typically requires 100 NFEs, leading to substantial latency ($\sim$40s on a single A100 for $\times1024$ resolution). Recent works distill to cut NFEs while preserving compositionality and aesthetics: LCM-style distillation for latent models (Luo et al., 2023), large-scale text-to-image distillation pipelines (e.g., PixArt-Delta (Chen et al., 2024b), SDXL distillation (Lin et al., 2024), FLUX-schnell (Labs, 2024)), and hybrid frameworks such as SANA-Sprint (Chen et al., 2025c) that combine teacher guidance with adversarial signals. Recently, Hunyuan-Image-2.1 (Team, 2025) explore MeanFlow (Geng et al., 2025) for mid-step acceleration (16 NFEs). Nevertheless, when targeting 1-2 NFEs on $\times1024$ text-to-image with 10B-20B backbones, these pipelines face practical barriers: dependence on frozen teachers, extra discriminators or score networks, unstable adversarial training, and prohibitive memory costs that hinder straightforward scaling.
Concurrent progress in few-step generative models.
Recent research has significantly diversified the landscape of efficient generative models, introducing novel training objectives and distillation paradigms. FACM (Peng et al., 2025) stabilizes consistency training by explicitly anchoring the shortcut objective to the instantaneous velocity field of a flow matching task. SoFlow (Luo et al., 2025) proposes a solution consistency loss that bypasses expensive JVP calculations, enabling efficient one-step generation. Adopting a policy-optimization perspective, $\pi$-Flow (Chen et al., 2025a) decouples the number of network evaluations from ODE integration steps via an imitation learning framework termed $\pi$-ID. Furthermore, Tong et al. (2025) introduces a data-free distillation paradigm that samples from the prior distribution to eliminate teacher-data mismatch, while Self-E (Yu et al., 2025) employs a self-evaluation mechanism to refine generation quality by leveraging the model’s own estimates. T2I-Distill (Pu et al., 2025) systematically evaluates sCM (Lu and Song, 2024) and MeanFlow (Geng et al., 2025) methods and developed efficient JVP kernels, offering practical guidelines for adapting them to open-domain text-to-image generation.
In the realm of methods with distribution matching, Decoupled-DMD (Liu et al., 2025a) and DMD-R (Jiang et al., 2025) offer new insights into the DMD objective; the former identifies CFG augmentation as the primary driver of acceleration while treating distribution matching as a regularizer, whereas the latter integrates reinforcement learning to enhance mode coverage during distillation. Targeting large-scale video generation, rCM (Zheng et al., 2025) combines continuous-time consistency models with score distillation regularization to resolve fine-detail degradation, and Transition Matching Distillation (TMD) (Nie et al., 2026) utilizes a decoupled architecture with a recurrent flow head to efficiently distill video diffusion models into few-step generators.
@@ B.1 Detailed Implementation on Multimodal Generative Models
For training OpenUni (Wu et al., 2025c) and Qwen-Image (Wu et al., 2025a), we used the same datasets as in our text-to-image experiments but excluded ShareGPT-4o-Image (Chen et al., 2025d). For exploration on full-parameter training on Qwen-Image, since DMD (Yin et al., 2024b) employs a regression loss, we need to pre-generate offline data using Qwen-Image. To ensure a fair comparison with established baselines, we randomly sample 50K prompts from the datasets we used, and generate images using Qwen-Image. For Qwen-Image-Edit (Wu et al., 2025a), we used only the part of editing split of the ShareGPT-4o-Image dataset.
@@ B.2 Detailed Implementation on Dedicated Text-to-image Models
Considering training datasets, we use BLIP-3o-60K (Chen et al., 2025b), Echo-4o (w/o multi-reference split) (Ye et al., 2025), and ShareGPT-4o-Image (Chen et al., 2025d). Together, these three instruction tuning datasets comprise approximately 200,000 text-to-image samples.
@@ B.3 Exploration on Image Editing
We conducted a preliminary exploration of our method on image editing tasks using a subset of approximately 15,000 square images from Chen et al. (2025d). We fine-tuned the Qwen-Image-Edit model (Wu et al., 2025a) using LoRA at a fixed $\times512$ resolution, with training configurations detailed in Table 6. Note that we use a low and fixed resolution during training; during testing, we use a resolution that is the same as the input image size. Our Qwen-Image-Edit-TwinFlow can effectively edit images in 2 to 4 NFEs. With 2 NFEs, it achieves a score of 3.47 on the ImgEdit, surpassing all multi-step models except Qwen-Image-Edit itself.